User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

Meeting 18th August

Chair: Liz

Present: Tiege, Jonny, Ross, Mas, Patrick, Bryn, Jamie, Lucia

Apologies: Lucas

Absent: Arwin, Jesse

General Issues:

  • Ho camera doesn't slew to the stuff in the drop down menu. Again broken. Don't scroll too excitedly. Someone can manually move it to fix. A persistent offset in x.
  • Email re. data transfer rd experiments not done. Warren!
  • Mid-sem break proposed meeting for Arwin. But too many people will be away so maybe not. Stay with how it is. But for Arwin and others who can't make meeting: update meeting notes pre meeting with any issues they want discussed.

Handover Issues:

  • Ho26m * aua024 illegal sample rate? can fix by loading the procedure file for aua022. Should we really be doing this though… BUT do it after schedule=!!! Yes, correct thing to fix. But should be fixed in the scheduling stage not observing anyway. Strange formatter thing…
  • 20K not reporting properly started on crds90… Nothing was on fire. But is ongoing and potentially the diode is having issues and giving non real values.Not high priority fix so keep an eye on it.
  • Modules? on both aua024 and crds90 Warren put a module in the wrong place. Is okay. Confusion is a problem but should be better when we make full transition to new stuff.
  • rxp reports wrong values again.Also on live page. If you change receivers it works, and panel is fine. But drive pc issue where it doesn't get correct information. Hopefully a reboot will fix, will do asap.
  • Yg * R4804 Mas? Antenna stuck? Tried to fix, and onsite person tried to fix. But just decided to work when it wanted. But Yg people have put in a temporary fix on Monday which may make the brake problem better. No problems were reported for r1 so can guess it is fixed (for now).
  • R4804 Network connection to maser. This is all okay? Vnc had an issue we thing. Maser seemed fine. Seems that a bad connection reports as battery power or something.Follow up with Brett is forthcoming.
  • Ke * R1804 drives stuck but fixed by reboot. Is this important? probably a power glitch but check that antenna=open. And reset drives in HMI but didn't work just from antenna=open. Seems okay…
/home/www/auscope/opswiki/data/pages/operations/agenda2017_08_03.txt · Last modified: 2017/08/17 03:32 by Elizabeth Iles